OpenStudy (anonymous):

how would i solve 4b-8-b=10b-3b? 3b-8=7b. 3b=15b?

OpenStudy (anonymous):

\[4b-8-b=10b-3b\]\[3b-8=7b\]-3b -3\[-8=4b\]/4 /4\[-2=b\]

OpenStudy (anonymous):

You wouldn't go from 3b-8=7b --> 3b=15b. It would be 3b - 8 = 7b -3b -3b -8 = 4b Then divide. /4 /4 b = -2

OpenStudy (anonymous):

oh ok i see what you mean thanks guys

OpenStudy (anonymous):

if i had a problem with fractions how would that work out? would i add to make one side a whole number

OpenStudy (anonymous):

Can you give us the problem?

OpenStudy (anonymous):

There is no need to mess with fractions in this problem.

OpenStudy (anonymous):

this is for a different problem. - 3/8y=-1/16. do i just eliminate one side?

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

\[-3/8y=-1/16\]\[(8y)(16)(-3/8y)=(-1/16)(8y)(16)\]\[(16)(-3)=(-1)(8y)\]\[-48=-8y\]/-8 /-8\[6=y\]

OpenStudy (anonymous):

i can see how you solved it but i dont understand the second line quite well. how 8y and 16 play in that part

OpenStudy (anonymous):

Use "The Granny Rule: Grannies are fair, what they give to one child they give to the other." Same in algebra, what you do to one side of the equation you do to the other. I multiplied both sides by 16 and 8y. On the left side the 8y's canceled out, on the right the 16's canceled out.

OpenStudy (anonymous):

oh ok i got it, i think thanks